OpenStudy (owlet):

Simplify: \(\sf sin(tan^{-1}x)\)

OpenStudy (owlet):

answer is \(\sf \large \frac{x}{1+x^2}\)

OpenStudy (owlet):

idk how they get it

OpenStudy (owlet):

correction, answer is \(\sf \large \frac{x}{\sqrt{1+x^2}}\)

OpenStudy (owlet):

@triciaal @Loser66 @Nnesha

Nnesha (nnesha):

we can draw a right triangle tan = opposite over adjacent so arctan x is same as x/1 \[\huge\rm \tan \theta =\frac{ opposite }{ adjacent }=\frac{ x }{ 1 }\]

Nnesha (nnesha):

 apply the Pythagorean theorem to find 3rd side of right triangle

OpenStudy (owlet):

thanks! got it :)

Nnesha (nnesha):

np :=0

Nnesha (nnesha):

:=)*